# BadPromptFL: A Novel Backdoor Threat to Prompt-based Federated Learning in Multimodal Models

### v1

@@ VI Experiments
TABLE I: Clean classification accuracy (ACC, %) and attack success rate (ASR, %) on different datasets under Non-IID client data distribution and varying prompt adaptation settings. The first row reports the 0-shot accuracy of the original model without tuning, the second row shows the 6-shot accuracy without any attack, and the third row presents the 6-shot accuracy and ASR when backdoor prompts are injected.
@@
RN50 86.00 - 17.20 - 55.60 - 61.40 - 40.40 - 66.00 - 37.50 - 40.40 -
@@
(a) RN50: Comparison of ACC and ASR across datasets.
(b) ViT: Comparison of ACC and ASR across datasets.
Figure 3: ACC and ASR performance of BadPromptFL compared to standard 0-shot and 6-shot FL across eight downstream datasets. The top row shows results on RN50, while the bottom row shows ViT. For each dataset, clean test accuracy (ACC) and attack success rate (ASR) are reported. BadPromptFL consistently achieves high ASR with minimal compromise in model utility.
@@
TABLE II: Clean classification accuracy (ACC, %) and attack success rate (ASR, %) of different defense methods evaluated independently on five natural image datasets. The results show the performance of contextual prompts on clean and backdoor tasks under different aggregation strategies, illustrating the impact of defense mechanisms on model robustness and utility. We select five datasets for defense evaluation to balance computational cost and ensure domain diversity, covering generic objects, fine-grained categories, scenes, actions, and textures.
@@
VI-A Experimental Settings
VI-A1 Model.
VI-A2 Datasets.
VI-A3 Federated Learning Setup.
We follow a full participation protocol in which all clients are involved in every communication round without client sampling. Federated training is conducted for 100 communication rounds. In each round, every client performs 5 local epochs of prompt optimization on its private data before uploading the updated prompt to the server. For the few-shot experiments, each client constructs its local training set by sampling 6 labeled examples per class (6-shot). To simulate realistic federated environments, we adopt a label-heterogeneous (non-IID) partitioning strategy, where each client holds data from a disjoint subset of classes. After aggregation, the updated global prompt is broadcasted to all clients and used to initialize the next round. All experiments use the same hyperparameters unless otherwise specified.
VI-A4 Backdoor Defenses.
  • Differential Privacy (DP) [54]: DP-SGD with noise multipliers of 0.02 and 0.2 applied to client prompt gradients.
VI-A5 Evaluation Metrics.
VI-B Experiments Results
VI-B1 Performance.
We evaluate the effectiveness of BadPromptFL in terms of clean classification accuracy (ACC) and attack success rate (ASR) across eight diverse datasets. Table I presents detailed results under three prompt tuning settings: 0-shot (no tuning), 6-shot (clean tuning), and 6-shot with backdoor injection.
Figure 4 provides a visual comparison of ACC and ASR under the three settings for both RN50 and ViT backbones. The top row (a) shows results for RN50, and the bottom row (b) corresponds to ViT. Across most datasets, BadPromptFL achieves a strong attack success rate with minimal degradation in classification performance, demonstrating its stealth and effectiveness in multimodal prompt-based federated learning.
VI-B2 Defense Evaluation.
Table II presents the clean accuracy (ACC) and attack success rate (ASR) achieved by different defense strategies evaluated independently on each dataset. Without any defense (FedAvg), the model exhibits consistently high ASR across datasets, confirming the vulnerability of prompt aggregation to backdoor injection.
Among the defense methods, Differential Privacy with a strong noise multiplier (DP0.2) is the most effective in reducing ASR, often approaching zero. However, this comes at the cost of substantial degradation in clean accuracy, dropping to 7.9% on Oxford Flowers and 23.8% on EuroSAT. MKrum and Foolsgold reduce ASR moderately while maintaining higher accuracy, but ASR remains above 60% on several datasets. Multi Metric achieves partial mitigation but still shows high ASR in most cases.
Overall, these results highlight a clear trade-off between backdoor robustness and clean performance. While stronger defenses such as Differential Privacy can substantially suppress ASR, this often comes at the expense of severe accuracy degradation. For example, DP0.2 reduces ASR close to zero but also lowers clean accuracy below the 0-shot baseline on several datasets, indicating that the model’s adapted prompts are effectively rendered unusable. MKrum and Foolsgold provide partial mitigation but still fail to reduce ASR to acceptable levels without significant performance loss. These observations suggest that existing defense strategies, primarily designed for gradient or weight-based poisoning, are insufficient in this context. Developing methods that can simultaneously preserve clean accuracy and resist prompt-level backdoor injection remains an important direction for future research.
@@ VII Conclusions
In this work, we present BadPromptFL, a novel backdoor attack that targets the emerging paradigm of prompt-based federated learning (PromptFL). Unlike traditional model-centric federated attacks, BadPromptFL exploits the unique vulnerability introduced by aggregating prompts—task-dependent embedding vectors that directly influence multimodal model behavior. By jointly optimizing visual triggers and poisoned prompt embeddings, adversarial clients can implant persistent backdoors without modifying the underlying model architecture or accessing global data.

### v3

@@ VI Experiments
TABLE I: Clean classification accuracy (ACC, %) and attack success rate (ASR, %) on different datasets under IID client data distribution and varying prompt adaptation settings. The first row reports the 0-shot accuracy of the original model without tuning, the second row shows the 6-shot accuracy without any attack, and the third row presents the 6-shot accuracy and ASR when backdoor prompts are injected.
@@
RN50 86.00 - 17.20 - 55.60 - 85.80 - 66.00 - 37.50 - 61.40 - 40.40 -
92.70 - 32.61 - 73.70 - 89.18 - 95.70 - 84.98 - 79.17 - 68.44 -
92.41 89.45 31.08 87.40 73.95 91.46 87.60 63.04 95.41 81.32 85.99 93.12 79.33 93.12 67.14 83.04
ViT 92.90 - 24.70 - 65.30 - 89.20 - 71.30 - 47.50 - 66.80 - 43.10 -
95.94 - 42.78 - 82.15 - 93.68 - 97.85 - 82.27 - 85.62 - 74.17 -
95.94 92.17 43.29 85.54 82.27 83.36 92.97 63.26 97.12 83.48 82.26 97.49 85.54 90.35 73.05 67.38
TABLE II: Clean classification accuracy (ACC, %) and attack success rate (ASR, %) on different datasets under Non-IID client data distribution and varying prompt adaptation settings. The first row reports the 0-shot accuracy of the original model without tuning, the second row shows the 6-shot accuracy without any attack, and the third row presents the 6-shot accuracy and ASR when backdoor prompts are injected.
ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR ACC ASR
RN50 86.00 - 17.20 - 55.60 - 85.80 - 66.00 - 37.50 - 61.40 - 40.40 -
@@
(a) RN50 (IID).
(b) ViT (IID).
(c) RN50 (Non-IID).
(d) ViT (Non-IID).
Figure 3: ACC and ASR performance of BadPromptFL compared to standard 0-shot and 6-shot FL across eight downstream datasets. The first row presents results under the IID setting for (a) RN50 and (b) ViT, while the second row shows results under the Non-IID setting for (c) RN50 and (d) ViT.
(a) Communication-round evolution of ACC and ASR for BadPromptFL compared to clean PromptFL on four representative datasets with RN50 (top row) and ViT (bottom row) backbones.
(b) Additional results on ACC and ASR evolution for four further datasets (Caltech101, FGVCAircraft, Oxford Pets, Stanford Cars), demonstrating consistent trends across domains.
Figure 4: Evolution of clean accuracy (ACC) and attack success rate (ASR) over 100 communication rounds for BadPromptFL compared to clean PromptFL. Subfigure (a) presents results on four representative datasets, while subfigure (b) shows supplementary experiments on additional datasets. Across both sets of experiments and architectures (RN50 and ViT), BadPromptFL rapidly attains high ASR in the early rounds while maintaining ACC close to the clean baseline throughout training, highlighting the effectiveness and stealthiness of the proposed attack.
VI-A Experimental Settings
VI-A1 Model
VI-A2 Datasets
VI-A3 Federated Learning Setup
We adopt a full-participation protocol in which all clients engage in every communication round without client sampling. Federated training is conducted for 100 communication rounds. In each round, every client performs 5 local epochs of prompt optimization on its private data before uploading the updated prompt to the server. For the few-shot experiments, each client constructs its local training set by sampling 6 labeled examples per class (6-shot). To simulate realistic federated environments, we evaluate our approach under both independent and identically distributed (IID) and label-heterogeneous (Non-IID) partitioning strategies, where in the Non-IID setting each client holds data from a disjoint subset of classes. After aggregation, the updated global prompt is broadcast to all clients and used to initialize the next round. Unless otherwise specified, all experiments are conducted under the same hyperparameter settings, with the learning rate for both clean and backdoor tasks set to 0.001 by default, and the trigger perturbation $\tau$ constrained within the range [8/255,8/255][-8/255,8/255].
VI-A4 FL Backdoor Defenses
  • Differential Privacy [38]: DP-SGD with noise multipliers of 0.02 applied to client prompt gradients.
VI-A5 Evaluation Metrics
VI-B Experiments Results
VI-B1 Main-task Accuracy and Attack Success Rate
We evaluate the effectiveness of BadPromptFL in terms of ACC and ASR across eight diverse datasets. Table I and Table II presents detailed results under three prompt tuning settings: 0-shot (no tuning), 6-shot (clean tuning), and 6-shot with backdoor injection.
Fig. 3 provides a visual comparison of ACC and ASR under three settings for both RN50 and ViT backbones. The top row (a) reports results under the IID setting across eight datasets, while the bottom row (b) corresponds to the Non-IID setting. Across most datasets, BadPromptFL achieves a consistently high attack success rate with minimal degradation in classification performance, demonstrating its stealth and effectiveness in multimodal prompt-based federated learning.
VI-B2 Dynamics over Communication Rounds
To further analyze the temporal behavior of BadPromptFL, we track the evolution of ACC and ASR over 100 communication rounds on four representative datasets. As shown in Fig. 4, BadPromptFL quickly achieves a high ASR within the first 10–20 rounds, indicating that the malicious prompts are effectively injected and aggregated early in the training process.
Despite the rapid escalation of ASR, the ACC remains close to that of the clean PromptFL baseline throughout the entire training, demonstrating the stealthiness of the attack. This stability in ACC is particularly evident in datasets such as Oxford Flowers and UCF101, where performance degradation is minimal despite ASR approaching 100%.
These results highlight two important characteristics of BadPromptFL in federated prompt learning: (1) high backdoor effectiveness can be achieved early, reducing the need for prolonged malicious participation, and (2) the main-task performance remains largely unaffected, making detection by standard performance monitoring highly challenging.
VI-B3 Effectiveness with Different Numbers of Malicious Clients
We further investigate the relationship between the proportion of malicious clients and the performance of BadPromptFL in terms of both ACC and ASR. Fig. 5 reports results under varying malicious-client ratios $\rho\in\{10\%,20\%,30\%,40\%\}$ across four representative downstream datasets (Oxford Flowers, EuroSAT, DTD, and UCF101) and two CLIP-based architectures (RN50 and ViT), evaluated under the non-IID setting.
[…]
These results highlight two key properties of BadPromptFL: (i) the attack can be effectively launched without requiring full control over the most clients, making it feasible in realistic threat models, and (ii) the conditional trigger mechanism enables sustained stealthiness and model performance regardless of the adversary’s scale of participation.
(a) RN50
(b) ViT
Figure 6: Attack performance under label heterogeneity with different model architectures: (a) RN50 and (b) ViT. Results are reported on four datasets with varying Dirichlet parameters $\alpha\in\{0.1,0.3,0.5,0.7,0.9\}$, showing that the attack success rate (ASR) remains consistently high while accuracy (ACC) only slightly declines under severe non-IID conditions.
Figure 7: Visualization of label heterogeneity across 20 clients under different Dirichlet distribution parameters. From left to right, the figures correspond to $\alpha=0.1$, $\alpha=0.5$, and $\alpha=0.9$. For all datasets except EuroSAT, we select 20 representative classes for visualization. In each figure, the vertical axis denotes clients, and each dot represents a class owned by a client. The dot size reflects the number of classes possessed by that client (larger dots indicate more classes), while the color intensity indicates the number of clients sharing the same class.
VI-B4 Effectiveness under Label Heterogeneity
To further investigate the robustness of our proposed attack in realistic federated learning scenarios, we evaluate its effectiveness under label heterogeneity, where each client only possesses data from a limited subset of classes. Fig. 6 reports the attack performance across four datasets (Oxford Flowers, EuroSAT, DTD, and UCF101) with varying degrees of heterogeneity, parameterized by $\alpha\in\{0.1,0.3,0.5,0.7,0.9\}$ in the Dirichlet distribution.
The results reveal that the ASR remains consistently high across different levels of heterogeneity, demonstrating that our backdoor design can effectively adapt to heterogeneous data distributions. In particular, even at the most skewed case ($\alpha=0.1$), the ASR exceeds 70% on all datasets, while the model utility measured by ACC only experiences a marginal decline. As $\alpha$ increases (i.e., the client data becomes more balanced), the ASR remains stable, and the ACC gradually improves, indicating that heterogeneity primarily affects the main-task performance rather than the backdoor activation.
These findings confirm that label heterogeneity does not significantly impair the attack’s effectiveness, highlighting the practicality and stealthiness of our method in realistic federated settings where non-i.i.d. data distributions are inevitable.
To better illustrate the impact of label heterogeneity, we visualize the client-class distributions on the EuroSAT dataset under three representative settings of $\alpha\in\{0.1,0.5,0.9\}$, as shown in Fig. 7. In this visualization, each dot corresponds to a client-class pair, where the dot size denotes the number of categories owned by the client, and the dot color reflects how many clients share the same class (darker colors indicate broader coverage).
From the figure, we observe that with a small $\alpha$ (e.g., $\alpha=0.1$), the distributions are highly skewed: each client only holds a few classes (large variation in dot sizes), and many classes are concentrated in only a limited set of clients (darker clusters). As $\alpha$ increases (e.g., $\alpha=0.9$), the distribution becomes much more balanced, with most clients covering a larger number of classes and each class being shared across multiple clients. This visualization clearly demonstrates the transition from highly heterogeneous to nearly homogeneous distributions as $\alpha$ grows, providing an intuitive understanding of how the Dirichlet parameter controls data heterogeneity in federated learning.
VI-B5 Image Quality Evaluation via PSNR and SSIM
To further assess the impact of our proposed attack on visual fidelity, we evaluate image quality using two widely adopted metrics: Peak Signal-to-Noise Ratio (PSNR [56]) and Structural Similarity Index (SSIM [57]). As reported in Table III, the reconstructed images maintain consistently high values across different datasets and architectures, indicating that the perturbations introduced by our method incur only negligible visual distortion. For instance, on the OxfordFlowers dataset, PSNR remains above 35 dB and SSIM exceeds 0.90 for both RN50 and ViT backbones, suggesting that the generated samples are nearly indistinguishable from clean counterparts to human perception. Similarly, for more challenging datasets such as EuroSAT and UCF101, our method sustains PSNR values around 35 dB and SSIM above 0.80, further confirming the imperceptibility of the injected transformations. These results demonstrate that our attack preserves image realism while effectively embedding the backdoor, thereby ensuring stealthiness against human inspection and potential detection mechanisms based on perceptual quality.
TABLE III: PSNR and SSIM for different datasets and architectures.
Dataset Architecture PSNR SSIM
Oxford Flowers RN50 35.1172 0.9035
ViT 34.9527 0.9004
EuroSAT RN50 34.8317 0.8072
ViT 34.8231 0.8069
DTD RN50 34.9538 0.9334
ViT 34.9038 0.9327
UCF101 RN50 35.2906 0.8683
ViT 35.1329 0.8644
VI-B6 Visualization of Clean and Poisoned Samples
To qualitatively assess the effect of backdoor injection, we visualize matched clean–poisoned pairs from four benchmark datasets. The visualization in Fig. 8 is arranged as an $8\times 6$ panel, where each dataset occupies two consecutive rows (from top to bottom: Oxford Flowers, EuroSAT, DTD, and UCF101). Within each row, the six columns are grouped into three pairs; in each pair, the left image is a clean sample and the right image is its poisoned counterpart. Consequently, each dataset presents six clean–poisoned pairs in total across the two rows.
In this study, the backdoor is implemented by applying imperceptible noise perturbations across the entire image. Unlike localized patch-based triggers, such global perturbations are visually indistinguishable from clean inputs and do not alter the semantic content of the image. As illustrated in Fig. 8, the poisoned samples appear nearly identical to their clean counterparts to human observers. This design ensures high stealthiness, making the poisoned inputs extremely difficult to detect by manual inspection, thereby highlighting the necessity of rigorous and automated defense mechanisms.
Figure 8: Visualization of clean–poisoned pairs across four datasets. The figure is an $8\times 6$ panel with two consecutive rows per dataset (top to bottom: Oxford Flowers, EuroSAT, DTD, UCF101). Within each row, the six columns form three pairs (left: clean; right: poisoned), yielding six pairs per dataset across the two rows.
Figure 9: t-SNE visualization of clean and backdoor samples across multiple datasets. Clean images (benign clients) form class-consistent clusters, while backdoor images (poisoned clients) exhibit clear separation in the embedding space. Despite the triggers being visually subtle, their injected prompts dominate representation learning, leading to a distinct backdoor subspace with partial overlap caused by residual clean features.
Figure 10: STRIP defense evaluation on the proposed backdoor attack across four datasets and two model architectures. The results indicate that STRIP struggles to effectively distinguish backdoor samples from clean samples, demonstrating the stealthiness of the attack.
VI-B7 t-SNE Visualization
To provide a more intuitive understanding of the embedding manipulation induced by BadPromptFL, we conduct t-SNE visualizations across multiple datasets. In our setup, we randomly sample clean images from several categories and generate corresponding backdoor images by injecting visually subtle triggers. This allows us to examine the distribution of prompt embeddings for both benign and poisoned samples under different aggregation scenarios.
As shown in Fig. 9, the embeddings of clean images from benign clients form compact and well-separated clusters that align with their semantic categories, reflecting consistent representation learning in standard federated prompt tuning. In contrast, once backdoor samples are introduced, their embeddings deviate from these clusters and form distinguishable subspaces in the global representation space. This separation demonstrates that poisoned prompts successfully carve out a latent decision boundary that facilitates backdoor activation at inference. Notably, despite the triggers being designed to be imperceptible, their semantic impact on the learned embeddings is sufficient to drive a clear divergence from the benign distribution.
Interestingly, we also observe that certain backdoor samples remain partially clustered with their clean counterparts. This phenomenon can be attributed to the residual contribution of original image features, which still influence the embedding process despite the presence of triggers. Such overlap indicates that the backdoor signal does not entirely erase the semantic information of the base image but instead coexists with it, thereby enhancing the stealthiness of the attack.
Overall, these visualizations highlight a critical insight: BadPromptFL enables poisoned prompts to embed a hidden yet persistent backdoor subspace into the global model. This subspace remains indistinguishable from benign clusters to human inspection but can be reliably exploited during inference. The coexistence of semantic alignment and malicious separation underscores both the robustness and stealthiness of our attack in federated multimodal learning.
@@ VI-B6 Visualization of Clean and Poisoned Samples
TABLE IV: Clean classification accuracy (ACC, %) and attack success rate (ASR, %) of different defense methods evaluated independently on five natural image datasets. The results show the performance of contextual prompts on clean and backdoor tasks under different aggregation strategies, illustrating the impact of defense mechanisms on model robustness and utility. We select five datasets for defense evaluation to balance computational cost and ensure domain diversity, covering generic objects, fine-grained categories, scenes, actions, and textures.
@@
VI-C Resistance to Backdoor Defense Methods
VI-C1 Resistance to FL Defenses
Table IV summarizes the ACC and ASR of BadPromptFL under various backdoor defenses across five natural image datasets for both RN50 and ViT architectures. Without defenses (FedAvg), BadPromptFL achieves consistently high ASR (often exceeding $90\%$) while maintaining competitive ACC, confirming the susceptibility of prompt aggregation to backdoor injection. Among the evaluated methods, Differential Privacy with a mild noise multiplier (DP) provides partial ASR reduction (e.g., from $92.68\%$ to $67.59\%$ on RN50–UCF101), but is insufficient to suppress attacks to acceptable levels and can significantly degrade ACC in some datasets (e.g., Oxford Flowers). Aggregation-based defenses such as MKrum and Foolsgold demonstrate dataset-dependent effectiveness: MKrum moderately reduces ASR while preserving higher ACC, yet still leaves ASR above $80\%$ in most cases; Foolsgold can drastically suppress ASR in certain settings (e.g., $1.69\%$ on RN50–UCF101) but fails in others (e.g., $77.77\%$ on RN50–Caltech101). The Multi Metric defense achieves strong suppression on specific dataset–model pairs (e.g., $2.07\%$ ASR on ViT–DTD) but suffers from very high ASR in other cases (e.g., $99.93\%$ on RN50–EuroSAT). These results indicate that while existing defenses can sometimes reduce prompt-level backdoor impact, none consistently balance robustness and clean utility, underscoring the need for defense mechanisms specifically tailored to prompt-based federated learning.
VI-C2 Resistance to STRIP Defense
To further evaluate the robustness of our proposed BadPromptFL against existing backdoor defenses, we adopt the STRIP [58], which is one of the most representative run-time detection techniques for poisoned inputs. The core idea of STRIP is to apply strong perturbations (e.g., superimposed images) to the incoming input and measure the entropy of the model’s predictions. For benign samples, the prediction distribution tends to fluctuate significantly under perturbations, leading to higher entropy. In contrast, poisoned samples usually maintain a stable prediction towards the target label, resulting in lower entropy values. This discrepancy enables STRIP to identify suspicious inputs.
Fig. 10 presents the normalized entropy distributions of benign and poisonous samples across different datasets and model architectures. Ideally, a clear separation between benign and poisoned samples would be observed, which facilitates reliable detection. However, the results reveal that in the presence of BadPromptFL, the entropy distributions of poisonous samples largely overlap with those of benign samples. Specifically, the distributions exhibit similar patterns without significant divergence, making it challenging for STRIP to set a robust threshold that balances false positives and false negatives.
These findings indicate that BadPromptFL can effectively bypass STRIP detection, since the poisoned samples are carefully crafted to mimic the entropy characteristics of benign ones. Consequently, the defense is unable to achieve reliable discrimination, highlighting the stealthiness and resilience of our attack against run-time detection mechanisms.
Figure 11: Effectiveness of CLP defense against BadPromptFL. We report clean accuracy (ACC) and attack success rate (ASR) across four datasets under IID and Non-IID settings with different purification coefficients $u$. The green dashed line indicates the zero-shot CLIP performance. Both ACC and ASR remain stable at large $u$, but sharply decline once $u$ falls below a threshold, illustrating the entanglement between benign and malicious features.
VI-C3 Resistance to CLP defense
To examine the effectiveness of the CLP defense against our attack, we evaluate BadPromptFL across four datasets under both IID and Non-IID settings while varying the purification coefficient $u$. As shown in Fig. 11, the green dashed line represents the zero-shot performance of the CLIP model as a reference. When $u$ is relatively large, both ACC and ASR remain stable, indicating that the purification has little effect on either normal or backdoored behavior. As $u$ decreases, ACC and ASR initially remain at high levels, but once $u$ drops below a certain threshold, both metrics decline in tandem. This pattern reveals that neurons responsible for benign features and those associated with backdoor features are strongly entangled, making it difficult for CLP to selectively eliminate malicious representations without also impairing clean performance. Although very small values of $u$ can suppress ASR, they simultaneously cause a severe degradation of ACC, effectively rendering the model unusable. These results demonstrate the inherent limitation of CLP in defending against prompt-based backdoor attacks, as purification cannot effectively decouple benign and malicious features while maintaining model utility.
VI-C4 Resistance to Neural Cleanse
Figure 12: Abnormal index values computed by Neural Cleanse for DTD and EuroSAT datasets using RN50 and ViT backbones. For DTD, the abnormal indices are 1.576 (RN50) and 1.075 (ViT); for EuroSAT, they are 1.448 (RN50) and 1.712 (ViT). Values below the typical anomaly threshold of 2 indicate that Neural Cleanse does not detect any backdoor in the BadPromptFL models.
In our experiments, Neural Cleanse (NC) was applied using 500 local images to perform trigger optimization. As illustrated in Fig. 12, all computed abnormal index values are below 2, indicating that NC fails to detect the backdoor in BadPromptFL. While a relatively large number of local samples were used for NC in this evaluation, such a setting does not conform to the federated learning paradigm, where the server lacks direct access to client data and each client possesses only a limited number of samples per class. These results demonstrate that, even under favorable conditions, NC is unable to detect the backdoor implanted by BadPromptFL.
@@ VII Conclusions
In this work, we present BadPromptFL, a novel backdoor attack that targets the emerging paradigm of prompt-based federated learning (PromptFL). Unlike traditional model-centric federated attacks, BadPromptFL exploits the unique vulnerability introduced by aggregating prompts—task-dependent embedding vectors that directly 55influence multimodal model behavior. By jointly optimizing visual triggers and poisoned prompt embeddings, adversarial clients can implant persistent backdoors without modifying the underlying model architecture or accessing global data.
@@
Ethical Statements
This work investigates BadPromptFL, a prompt-based backdoor attack framework in federated learning, with the sole purpose of uncovering and understanding potential security vulnerabilities in prompt aggregation for multimodal models such as CLIP. Our experiments are conducted in controlled environments without deployment in real-world systems, and no harmful models or triggers are released. The intention of this research is purely defensive: by systematically analyzing the weaknesses of current FL settings, we aim to provide the community with clear evidence of existing risks, thereby encouraging the design of more robust and secure learning frameworks. We firmly oppose any malicious use of the techniques described herein and stress that our findings should only serve as a warning and guidance for strengthening the safety of AI systems in practical deployments.